Tag Archives: section staining

Tuesday 03/05/13

9:50 A – 7:30 P, 8:30 P – 10:50 P General Discussing replies for Banff invites. Student substitutions add to bottom of list Still need to write more acknowledgement of replies. STORM Analysis launch analysis of all unprocessed datafiles. ESATA … Continue reading

Posted in Summaries | Tagged , , , , | Comments Off on Tuesday 03/05/13

Monday 03/04/13

9:45 A – 7:30 P, 9:00 P-11:30 P General stuff Group meeting: see notes send out invites to Banff workshop Continue working on writing referee review Fly work flipped all bottle stocks tossed all Mat alpha tub bottles currently expanding: … Continue reading

Posted in Summaries | Tagged , , | Comments Off on Monday 03/04/13

Saturday 03/02/13

10:00A – 7:00P, 8:30P-10:30P Writing stuff Revised Banff Invites letter sent to co-organizers write to van Steensel about R code Review/referee article read clean through manuscript wrote up all critiques. To do: write up manuscript strengths and contributions to field. … Continue reading

Posted in Summaries | Tagged , | Comments Off on Saturday 03/02/13

Friday 03/01/13

8:10 A – 8:10P, 9:30P – 11:00P Confocal helping scan Hazen’s sections Embryo section staining BXC DFISH NO BX-C staining detectable after post-hybe hot-wash (50-60C in waterbath, 25 min). Repeat staining, try short coldwash and then check cells for signal. … Continue reading

Posted in Summaries | Tagged , , | Comments Off on Friday 03/01/13

Thursday 02/28/13

10:20 A -6:45P, 8:15P – 11:55P (workout 8:30-10:20A) Probe Making (1P-6P @ Longwood) emPCR for 100kb yellow, green and black regions (at Wu lab) see new protocol 35-50 ng/uL yield (11 uL total volume) Notebook Looking in to version managing … Continue reading

Posted in Summaries | Tagged , , | Comments Off on Thursday 02/28/13

Wednesday 02/27/13

9:00 A – 10:00 P Embryo collection Flipped Esc cage — good looking overnight collection Probe design schematic of probes design and send to Brian new probe sequences for Custom Array order. working on new probe sets for colors project … Continue reading

Posted in Summaries | Tagged , , | Comments Off on Wednesday 02/27/13

Sunday 12/09/12

10:20 A – 7:30P, 9:30-11:00P STORM 750 1600 mA = 18 mW 1.45 mW 405 200 mW set 647 Made new Glox 750 not switching well. ┬áNot sure if it’s just low labeling (405 looked okay but not strong). 647 … Continue reading

Posted in Summaries | Tagged , , , , , | Comments Off on Sunday 12/09/12